### v1

# On the origin of exponential operator growth in Hilbert space

Abstract
By analysing the growth of Krylov complexity for some simple models, we find that the exponential growth of an operator in Hilbert space can be predicted from the off-diagonal decay of the operator matrix elements in the system eigenbasis. When the decay is exponential or slower, the Krylov complexity grows exponentially; when it is algebraic or slower, the growth rate is maximal. As a result, non-chaotic systems can generate exponential or even maximal operator growth, and chaotic systems can show sub-exponential growth; we give examples of both.
preprint: APS/123-QED
There has been much recent interest in the dynamics of operators in many-body quantum systems [1, 2, 3, 4], with particular focus on its connection with thermalisation, quantum chaos and ergodicity. It is already clear [5, 6] that there is no simple correspondence between exponential operator growth or ‘fast scrambling’ (as measured by out-of-time-ordered correlators (OTOC) [7, 8]) and ‘quantum chaos’ (as defined by energy-level statistics) . Various chaotic systems have been found to give exponential operator growth [9, 10, 11], but so have a number of simple integrable systems [5, 6].
Most recent work has been concerned with operator growth in Hilbert space [12, 13, 14], at temperatures that are sufficiently high that there are enough accessible states to give sustained growth until the space of available states is saturated. Under this formalism, when operator growth is exponential, the growth rate $2\alpha$ appears to be limited by the ‘universal’ (i.e. system-independent) bound
$\alpha(T)\leq{\pi k_{\text{B}}T}/{\hbar}$ (1)
This bound has the same value as an earlier-identified bound on OTOCs [8], but the two bounds are not equivalent, since it is easy to construct a few-body system that maximises $\alpha$ [13], whereas the OTOC bound is saturated only by certain many-body systems (e.g. a quantum model of a black hole [8] and the SYK model of strongly interacting majorana fermions [15]).
The purpose of this letter is to analyse exponential growth in a variety of systems under thermal equilibrium, and to show that it appears to be caused solely by the off-diagonal energy-dependence of the operator in the system eigenbasis. This property also determines whether the growth rate reaches the maximum bound given by Eq. (1). Using this diagnostic, it is easy to find examples of chaotic systems which do not give rise to exponential growth, and integrable systems which do, including integrable systems which give maximal growth.
To examine operator growth under conditions of thermal equilibrium, we follow refs. [12] in generating a Krylov basis from an operator $\hat{O}$, using the Lanczos sequence
${\mathcal{A}}_{n}=\hat{\mathcal{L}}\hat{O}_{n-1}-b_{n-1}\hat{O}_{n-2}$
$\hat{O}_{n}=b_{n}^{-1}{\mathcal{A}}_{n}$
$b_{n}=\sqrt{\langle{\mathcal{A}}_{n}|{\mathcal{A}}_{n}\rangle_{\beta}}$ (2)
where $\hat{\mathcal{L}}$ is the quantum Liouvillian superoperator
$\hat{\mathcal{L}}=[\hat{H}$$\cdot$$\hbar$ (3)
and
$\langle{\mathcal{A}}_{n}|{\mathcal{B}}_{n}\rangle_{\beta}=\frac{1}{Z}\text{Tr}[e^{-\beta\hat{H}/2}{\mathcal{A}}_{n}e^{-\beta\hat{H}/2}{\mathcal{B}}_{n}]$ (4)
The Krylov complexity (which monitors the growth of $\hat{O}$ in Hilbert space) is defined to be
$C_{O}^{K}(t)=\sum_{n=0}^{\infty}n|\varphi_{n}(t)|^{2}$ (5)
where the occupancy functions $\varphi_{n}(t)=i^{-n}\langle\hat{O}_{n}|\hat{O}(t)\rangle$ satisfy
$\frac{\partial\varphi_{n}(t)}{\partial t}=-b_{n+1}\>\varphi_{n+1}(t)+b_{n}\>\varphi_{n-1}(t)$ (6)
with $\varphi_{n}(0)=\delta_{n0}$. When
$C_{O}^{K}(t)\sim e^{2\alpha t}$ (7)
the operator is considered to grow exponentially in Hilbert space, and the exponent $\alpha$ is limited by Eq. (1). Equation (7) can be shown to be equivalent to linear growth
$b_{n}\sim\alpha n$ (8)
of the Lanczos coefficients $b_{n}$, and to exponential growth of the form
$\mu_{2n}\sim\left(\frac{4n\alpha}{e\pi}\right)^{2n}$ (9)
of the moments
$\mu_{2n}=\frac{1}{Z}{\partial^{2n}\over\partial t^{2n}}\left.\text{Tr}[e^{-\beta\hat{H}/2}\hat{O}(0)e^{-\beta\hat{H}/2}\hat{O}(t)]\right|_{t=0}$ (10)
of the time-correlation function defined as $C(t)=\langle\hat{O}(0)|\hat{O}(t)\rangle_{\beta}$.
[…]
Figure 1: Illustration of (a) power-law, (c) exponential and (e) Gaussian decay of off-diagonal matrix elements of an operator in the energy eigenbasis of a random-matrix ensemble. (b), (d) and (f) The corresponding Lanczos sequence averaged over the ensemble [16]. The red dashed line indicates the maximal growth as in (1).
In what follows, we will represent $\hat{O}$ in the basis set of eigenstates of $\hat{H}$. The off-diagonality of $\hat{O}$ in this basis is obviously crucial to its growth, since the Lanczos sequence vanishes entirely if $\hat{O}$ is diagonal. For a harmonic oscillator, the position operator $\hat{x}$ has matrix elements
$x_{lk}=\sqrt{\frac{\hbar}{2m\omega}}(\sqrt{k+1}\>\delta_{l,k+1}+\sqrt{k}\>\delta_{l,k-1})$ (11)
which cause the Lanczos sequence to terminate at $n=2$. But in general, the (symmetry-allowed) off-diagonal matrix elements of $\hat{O}$ are non-zero and typically decay as some function of the off-diagonal energy-difference $\omega$.
Let us now take an operator $\hat{O}$ in the eigenbasis of a ‘chaotic’ system which has energy-level spacings that satisfy Wigner-Dyson statistics and which obeys the Eigenstate Thermalisation Hypothesis (ETH) [17, 18]. Such an operator can be written
$O_{lk}:=\langle l|\hat{O}|k\rangle=O(\bar{E})\,\delta_{lk}+e^{-S(\bar{E})/2}\,f_{O}(\bar{E},\omega)\,R_{lk}$ (12)
[…]
$O_{lk}=\frac{1}{\sqrt{\mathcal{D}}}R_{lk}$ (13)
Expanding the moments $\mu_{2n}$ of Eq. (9) in the basis of energy eigenstates, we obtain
$\mu_{2n}=\frac{1}{Z}\sum_{l=1}^{\infty}\sum_{k=1}^{\infty}(E_{l}-E_{k})^{2n}e^{-\beta(E_{l}+E_{k})/2}|O_{lk}|^{2}$ (14)
We can rewrite this expression as
$\mu_{2n}$ $\approx\frac{\int_{0}^{\infty}d\bar{E}\>e^{S(\bar{E})-\beta\bar{E}}\int_{-2\bar{E}}^{2\bar{E}}d\omega\>\omega^{2n}|O(\bar{E},\omega)|^{2}}{\int_{0}^{\infty}d\bar{E}\>e^{S(\bar{E})-\beta\bar{E}}}$ (15)
provided the number of thermally accessible states is sufficiently high that we can replace the sum by an integral using [19]
$\sum_{l=1}^{\infty}\rightarrow\int dE_{l}\>e^{S(E_{l})},$ (16)
Averaging over the random variable $R_{lk}$, we obtain
$\langle\mu_{2n}\rangle$ $\sim\frac{\int_{0}^{\infty}d\bar{E}\>e^{S(\bar{E})-\beta\bar{E}}\int_{-2\bar{E}}^{2\bar{E}}d\omega\>\omega^{2n}|f_{O}(\bar{E},\omega)|^{2}}{\int_{0}^{\infty}d\bar{E}\>e^{S(\bar{E})-\beta\bar{E}}}$
$=2\int_{0}^{\infty}d\omega\>\omega^{2n}e^{-\beta|\omega|/2}$
$=\frac{2^{2n+2}(2n)!}{\beta^{2n+1}}$
$\sim\left(\frac{4n}{\beta e}\right)^{2n}$ (17)
which corresponds to exponential growth of $C_{O}^{K}(t)$ at the maximal rate $\alpha=\pi/\beta$ permitted by Eq. (1) (see Eq. (9)) setting $\hbar=1$. Numerical propagation of Eq. (2) confirms this prediction (indicating that Eq. (16) is a valid assumption at temperatures sufficiently high for exponential growth), with $b_{n}$ growing linearly at the maximal growth rate—see Fig. 1.
If we repeat the steps that led to Eq. (17), but with the structure function $f_{O}$ exponentially suppressed in $\omega$ as:
$f_{O}(\bar{E},\omega)\sim e^{-\gamma|\omega|}$ (18)
we obtain
$\mu_{2n}$ $\sim\frac{\int_{0}^{\infty}d\bar{E}\>e^{-\beta\bar{E}}\int_{-2\bar{E}}^{2\bar{E}}d\omega\>\omega^{2n}e^{-2\gamma|\omega|}}{\int_{0}^{\infty}d\bar{E}\>e^{-\beta\bar{E}}}$
$=\frac{2^{2n+2}(2n)!}{(\beta+4\gamma)^{2n+1}}$
$\sim\left[\frac{4n}{(\beta+4\gamma)e}\right]^{2n}$ (19)
i.e. $\hat{O}$ is still predicted to grow exponentially, but at a sub-maximal rate $\alpha=\pi/(\beta+4\gamma)$, which is again confirmed by the numerical results in Fig. 1. Such exponential (or Gaussian) suppression is a result of higher-order many-body processes required for transitions between states with widely separated energies [19, 20, 21, 22].
If we were now to retrace these steps for an $f_{O}$ which decays faster than exponentially as a function of $\omega$, it is clear that the expression we would derive for $\mu_{2n}$ would grow more slower than $\sim n^{2n}$, which would correspond to slower than exponential growth in $C_{O}^{K}(t)$ (and sub-linear growth with in $b_{n}$). This is motivated by the fact that quantum thermal fluctuations allow transitions between all pairs of states mediated by an operator $\hat{O}$ and enable its exponential growth as long as it is unhindered by the internal fluctuations [19] of operators. This thereby indicates a purely statistical origin of the bound on the growth rate similar to the bound on the quantum Lyapunov exponent [11, 23, 24, 25]. Numerical propagation of Eq. (2) also confirms this prediction—see Fig. 1.
These random-matrix results suggest a simple diagnostic for predicting whether an operator $\hat{O}$ will grow exponentially. The growth rate is determined by the dependence of $\hat{O}$ on the off-diagonal energy difference $\omega$ in the system eigenbasis. When $\hat{O}$ decays exponentially or more slowly with $\omega$, the Krylov complexity will grow exponentially, and when it decays algebraically or more slowly, this growth rate will maximise Eq. (1); otherwise the growth will be slower than exponential.
We now discuss a few examples which show that this simple picture appears to hold in more realistic models and is consistent with reported results in the literature.
First, consider the growth of the operator $\hat{\mu}=\hat{x}^{q}$ for a harmonic oscillator. For a large enough value of $q$ and a high enough temperature we can expect a period of sustained growth before the Lanczos series terminates at $n=q+1$. The off-diagonal matrix elements of $\hat{\mu}$ can be shown to decay as
$\langle l|\hat{\mu}|k\rangle\sim\exp\left(-\frac{|l-k|^{2}}{2q}\right),$ (20)
(see the SI). On the basis of the above, we therefore expect the operator growth to be slower than exponential, and this is confirmed numerically in Fig. 2 (using $m=1$ and $\beta=1$). Let us now define an operator $\hat{O}$ such that the matrix elements (in the harmonic-oscillator eigenbasis) take the form $\hat{O}_{lk}=R_{lk}$. Averaging over the $R_{lk}$, we obtain
$\langle\mu_{2n}\rangle$ $=\frac{1}{Z}\sum_{l=1}^{\infty}\sum_{m=1}^{\infty}(E_{l}-E_{m})^{2n}e^{-\beta(E_{l}+E_{m})/2}\langle|O_{lm}^{2}|\rangle$
$=\frac{1}{Z}\sum_{l=1}^{\infty}\sum_{m=1}^{\infty}[\hbar\omega(l-m)]^{2n}e^{-\beta\hbar\omega(l+m+1)/2}$
$=2(\hbar\omega)^{2n}\text{Li}_{-2n}(e^{-\beta\hbar\omega/2})$ (21)
where $\text{Li}_{-n}(z)$ is the negative polylogarithm function. As $\beta\hbar\omega\rightarrow 0$, we obtain
$\langle\mu_{2n}\rangle\rightarrow\frac{\omega^{2n}}{(\beta\omega/2)^{2n+1}}(2n)!\rightarrow\left(\frac{4n}{\beta e}\right)^{2n}$ (22)
In other words, even a harmonic oscillator can give rise to exponential operator growth that maximises Eq. (1) (in the high-temperature limit), provided the operator $\hat{O}$ decays sufficiently slowly as a function of $\omega$.
[…]
Figure 2: Plots of Lanczos sequence for (a) $\hat{O}=\hat{x}^{100}$ in a harmonic oscillator $V(x)=x^{2}/2$ (blue dots); the dashed line corresponds to a fictional operator with approximate matrix elements as in (20) indicating Gaussian decay of matrix elements. (b) $\hat{O}=\hat{x}$ in a quartic anharmonic oscillator $V(x)=x^{4}$ (yellow) - the artificial transition in the growth rate at $n\sim 25$ is due to the limit in numerical precision while computing position matrix elements; blue dots correspond to a fictional operator with (sub)exponentially decaying matrix elements as in (25). (c) $\hat{O}=\hat{x}$ in a one-dimensional infinite box (length $L=10$) at various temperatures. Inset: Plot of the growth rate against temperature indicating maximal growth.
One could argue that this last example is artificial, but more realistic examples of exponential or even maximal growth in non-chaotic systems can easily be found. For example, consider a one-dimensional system with potential
$V(x)=x^{p},$ (23)
where $p>2$ is an even integer. Using WKB theory [26, 27], the energy eigenvalues can be approximated by
$E_{k}\approx C(p)\>k^{2p/(p+2)}$ (24)
and the matrix elements of $\hat{x}$ by
$x_{lk}\approx\exp({-L_{lk}})\quad\quad|l-k|>>1$ (25)
where
$L_{lk}=J(p)|l-k|\sim|E_{l}^{(p+2)/2p}-E_{k}^{(p+2)/2p}|,$ (26)
and $C(p)$ and $J(p)$ are functions of $p$ (see the SI). For large |nm||n-m|,
$L_{lk}\sim|E_{l}-E_{k}|^{(p+2)/2p},$ (27)
and therefore $x_{lk}$ decays sub-exponentially with $\omega$ (since $(p+2)/2p\leq 1$). On the basis of the above, we therefore expect $O=\hat{x}$ to grow exponentially but not maximally. This prediction is confirmed numerically in Fig. 2, which was generated by propagating Eq. (2) (with $m=1$, $\beta=1$) with the $x_{lk}$ evaluated (i) using the numerically calculated eigenfunctions and (ii) analytically, using the semiclassical approximation of Eq. (25). Note the presence of a potentially misleading numerical artefact, in which the growth rate obtained in (i) appears to jump upwards at about $n=25$ to the maximal rate. This jump happens because the off-diagonal matrix elements for $\omega>75$ are dominated by machine errors and thus resemble random matrix elements (which do not decay with $\omega$).
As $p\rightarrow\infty$, the system of Eq. (23) tends to a particle in a one-dimensional box, for which
$x_{mn}=\begin{cases}\frac{8nm}{\pi^{2}(n^{2}-m^{2})^{2}}\;\;\;&m+n\text{ odd}\\ 0&\text{otherwise}\end{cases}$ (28)
Since these elements decay algebraically with $\omega$, we expect that $\hat{x}$ will grow at the maximum rate $\alpha=\pi/\beta$. Numerical propagation of Eq. (2) over a range of temperatures confirms this prediction over a wide temperature range—see Fig. 2.
Since a particle in a one-dimensional box gives maximal exponential growth, it follows immediately that a particle in a two-dimensional rectangular box also gives maximal growth (since the dynamics is equivalent to that of two decoupled one-dimensional particles). The growth of $\hat{x}$ for a particle in a rectangular box is plotted in Fig. 3, where it is shown to follow closely the growth reported in ref. [13] for a stadium billiard (until the growth stops at saturation). This shows that the maximal growth achieved by the latter is not a signature of quantum chaos, but rather a consequence of the algebraic decay of $x_{lk}$ with $\omega$, which results from confining the particle to a box.
[…]
Figure 3: (a) Lanczos sequence for $\hat{O}=\hat{x}$ in a stadium billiard (dots) with area 11 and the parameter length-radius ratio $a/R=1.0$ as in [13] at various temperatures. The dashed lines correspond to the Lanczos sequence for a rectangular box of similar dimensions as shown in the inset. (b) Lanczos sequence for the flip-flop operator $\hat{B}$ in (29) for the XXZ model with NN (green) and NNN (blue) coupling. The yellow dots corresponds to the operator truncated to an eigenbasis of size $N=2000$. Inset: Plots of the structure function at $E=0$ for each of the cases, showing exponential (NNN) and Gaussian (NN) decay.
As a final example, we consider the growth of the next-nearest-neighbour flip-flop operator
$\hat{B}=\frac{1}{L}\sum_{i=1}^{L}\hat{S}_{i}^{+}\hat{S}_{i+2}^{-}+\hat{S}_{i}^{-}\hat{S}_{i+2}^{+}$ (29)
for an XXZ Hamiltonian with nearest-neighbour (NN) and next-nearest-neighbour (NNN) coupling
$\begin{split}\hat{H}_{\text{XXZ}}=J_{1}\left(\sum_{i=1}^{L}\hat{S}_{i}^{x}\hat{S}_{i+1}^{x}+\hat{S}_{i}^{y}\hat{S}_{i+1}^{y}+\Delta_{1}\hat{S}_{i}^{z}\hat{S}_{i+1}^{z}\right)+\\ J_{2}\left(\sum_{i=1}^{L}\hat{S}_{i}^{x}\hat{S}_{i+2}^{x}+\hat{S}_{i}^{y}\hat{S}_{i+2}^{y}+\Delta_{2}\hat{S}_{i}^{z}\hat{S}_{i+2}^{z}\right)\end{split}$ (30)
where $i$ satisfies periodic boundary conditions. The parameter $J_{1}$ controls the nearest-neighbour (NN) coupling between sites, and $J_{2}$ controls the next-nearest-neighbour (NNN) coupling. We consider two different parameterisations of this model, an ‘NN’ model, in which $L=20$, $J_{1}=1.0$, $\Delta_{1}=0.55,\Delta_{2}=0.5$, and $J_{2}=0$, and an ‘NNN’ model which uses the same parameters except that $J_{2}=1$. The NN model is integrable and the matrix elements of $\hat{B}$ decay exponentially with $\omega$ for $N<2000$ and have Gaussian decay for $N>2000$; the NNN model is chaotic [28, 20] (with Wigner-Dyson statistics) and the matrix elements of $\hat{B}$ decay exponentially. We thus expect the NN model to give exponential growth initially, followed by slower growth, and the NNN model to give exponential growth. These predictions are confirmed by the numerical results in Fig. 3, generated by propagating Eq. (2) at $\beta=1$. To check that it is the $\omega$-dependence of $\hat{B}$ which produces the exponential growth in the NNN model (not the chaoticity), we rerun the NN calculations with the $\hat{B}$ matrix artificially truncated at $N=2000$ (so that all non-zero elements decay exponentially with $\omega$); as expected, the growth is purely exponential (Fig. 3).
In summary, it appears that whether an operator grows exponentially depends solely upon the $\omega$-dependence of its matrix elements in the system eigenbasis. Many chaotic systems will show exponential operator growth because the strong coupling between eigenstates that makes them chaotic will result in operator matrices which decay exponentially or more slowly with $\omega$. But this need not be the case: one can construct chaotic model systems for which the decay is faster and which therefore give operator growth that is slower than exponential. On the other hand, one can easily construct integral model systems in which the decay is exponential or slower and therefore results in exponential growth. In the latter case the strong coupling will be produced by some aspect of the system which is not related to chaos. For example, in the $V(x)=x^{p}$ case treated above, the WKB analysis [26, 27] shows that the slow decay of $\hat{x}$ with $\omega$ is caused by tunnelling at the outer turning point.
An obvious application of these findings is that it should never be necessary to propagate the Lanczos series in order to predict whether an operator $\hat{O}$ grows exponentially and, if so, whether the growth rate maximises the bound Eq. (1). Instead, these properties can be predicted immediately from the dependence of $\hat{O}$ on $\omega$. An understanding of operator growth is central to understanding concepts such as eigenstate thermalisation and scrambling in many body systems and this work relates it to properties of single-body systems, a connection which has been long sought after[29, 30]. Our theory also motivates the design of quantum quenches in anharmonic one-dimensional continuous quantum systems at finite temperatures to probe the nature of the resultant nonequilibrium state [31, 32].
VGS and SCA acknowledge funding from a Leverhulme Trust research grant. VGS, JMR and SCA contributed equally to this work.